# A Recipe For Obtaining Algebraic Addition Theorems Of The Weierstrass Elliptic Function

### v1

@@ 1 Introduction
Let $\Omega=(2\omega_{1},2\omega_{2})\in\mathbb{C}^{2}$ be such that $\text{Im}(\omega_{2}/\omega_{1})>0$ and $\Lambda_{\Omega}=2\omega_{1}\mathbb{Z}+2\omega_{2}\mathbb{Z}$. The Weierstrass elliptic function Weierstrass-$\wp$ is defined as
@@
The function Weierstrass-$\wp$ satisfies the well-known algebraic differential equation
$\left(\wp^{\prime}(z)\right)^{2}=4\wp(z)^{3}-g_{2}\wp(z)-g_{3}=4(\wp(z)-e_{1})(\wp(z)-e_{2})(\wp(z)-e_{3}).$ (1)
@@
$\begin{vmatrix}1&\wp(z_{0})&\wp^{\prime}(z_{0})&\cdots&\wp^{(n-1)}(z_{0})\\ 1&\wp(z_{1})&\wp^{\prime}(z_{1})&\cdots&\wp^{(n-1)}(z_{1})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&\wp(z_{n})&\wp^{\prime}(z_{n})&\cdots&\wp^{(n-1)}(z_{n})\end{vmatrix}=(-1)^{\frac{n(n-1)}{2}}1!2!\cdots n!\frac{\sigma(z_{0}+z_{1}+\cdots+z_{n})}{\sigma(z_{0})^{n+1}\cdots\sigma(z_{n})^{n+1}}\prod_{0\leq i<j\leq n}\sigma(z_{i}-z_{j}),$ (2)
@@
Frobenius-Stickelberger identities may be used to obtain algebraic addition theorems for the function Weierstrass-$\wp$ (for example, see [1, 8]). For more related determinant identities we refer the reader to [4, 5, 6, 7, 9]. The classical addition theorem for Weierstrass-$\wp$ is given as follows.
@@
For recent advances in addition theorems for elliptic functions, see [2, 3]. In this note, we discuss a recipe to obtain algebraic addition theorems for Weierstrass-$\wp$, expressing $\wp(z_{1}+\cdots+z_{\ell})$ in terms of $\wp^{(j)}(z_{i})$ for predetermined $j$.
@@ 2 Main Results
$\begin{vmatrix}\wp^{(k_{1})}(z_{1})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \wp^{(k_{1})}(z_{2})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \wp^{(k_{1})}(z_{\ell})&\wp^{(k_{2})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\end{vmatrix}\neq 0.$ (3)
@@
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{j})=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(z_{j})$ (4)
holds for every $j=1,\ldots,\ell$. Let us denote $z=z_{1}+\cdots+z_{\ell}$. Finally, assume that $z\notin\Lambda_{\Omega}$ and both of the equations $n_{a}=k_{b}$, $\gamma_{a}=\lambda_{b}$ do not hold at the same time.
We use Liouville’s theorems and elementary facts about elliptic functions frequently without mention. We first prove the determinant form of the addition theorem.
@@
By our assumptions, $\psi_{n,k,\gamma,\lambda}$ is an elliptic function of order $\max(n_{a},k_{b})+2=\ell+1$ with a pole of order $\ell+1$ at $s=0$. Since the equation (4) holds for any $j=1,\ldots,\ell$, the function $\psi_{n,k,\gamma,\lambda}$ has $\ell$ distinct zeros, namely $z_{1},\ldots,z_{\ell}$. Let $w$ denote the $\ell+1$th zero. Using Liouville’s theorem we obtain $w+\sum_{i=1}^{\ell}z_{i}\equiv 0\bmod\Lambda_{\Omega}$ and $w\equiv-z\bmod\Lambda_{\Omega}.$ Thus $-z$ is a root of $\psi_{n,k,\gamma,\lambda}$ and we get
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(-z).$ (5)
Putting equations (4) and (5) together gives
$\begin{pmatrix}\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\wp^{(k_{1})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\wp^{(k_{1})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\wp^{(k_{1})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)&\wp^{(k_{1})}(-z)&\cdots&\wp^{(k_{\ell})}(-z)\end{pmatrix}\begin{pmatrix}-1\\ \lambda_{1}\\ \vdots\\ \lambda_{\ell}\end{pmatrix}=-\begin{pmatrix}\psi_{n,k,\gamma,\lambda}(z_{1})\\ \psi_{n,k,\gamma,\lambda}(z_{2})\\ \vdots\\ \psi_{n,k,\gamma,\lambda}(z_{\ell})\\ \psi_{n,k,\gamma,\lambda}(-z)\end{pmatrix}=0.$ (6)
Since the vector on the left-hand side of (6) is non-zero, the matrix on the left-hand side is singular. Therefore we have
@@
We remark that since $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and Weierstrass-$\wp$,$\wp^{(i)}$ satisfy an algebraic relationship for every $i$, Theorem 2.1 is indeed an algebraic addition theorem.
@@
For all $i\in\mathbb{N}$, $\wp^{(2n)}\in\mathbb{Q}[g_{2},g_{2}][\wp]$ and $\wp^{(2n+1)}\in\wp^{\prime}\mathbb{Q}[g_{2},g_{2}][\wp]$. Furthermore $\wp^{(2n)}$ is a polynomial in Weierstrass-$\wp$ of degree $n+1$ and $\wp^{(2n+1)}/\wp^{\prime}$ a polynomial in Weierstrass-$\wp$ of degree $n$.
@@
The proof follows immediately after induction and equation (1). ∎
@@
$\varphi_{n,k,\gamma,\lambda}(s)=\left(\sum_{2\nmid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\nmid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}-\left(\sum_{2\mid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\mid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}$ (7)
@@
By Lemma 2.2 and equation (1), the first sum in (7) is in $\left(\wp^{\prime}\mathbb{C}[\wp]\right)^{2}$ and also in $\mathbb{C}[\wp]$. Similarly the second sum in (7) is a polynomial in $\mathbb{C}[\wp]$. Therefore $\varphi_{n,k,\gamma,\lambda}\in\mathbb{C}[\wp]$. Again by Lemma 2.2 and (1), the first sum is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{2\nmid n_{i},k_{j}}(n_{i},k_{j})+2$ and second sum is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{2\mid n_{i},k_{j}}(n_{i},k_{j})+2$. Thus $\varphi_{n,k,\gamma,\lambda}$ is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{i,j}(n_{i},k_{j})+2=\ell+1$. ∎
@@
Let $S_{r}(x_{1},\cdots x_{n})$, $k\leq n$ denote the elementary symmetric polynomials defined by
$S_{r}(x_{1},\ldots,x_{n})=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}x_{i_{1}}\ldots x_{i_{k}}.$
@@
Therefore $\wp(z_{1}),\ldots,\wp(z_{\ell})$ are distinct roots of $\varphi_{n,k,\gamma,\lambda}$. Using the fact that $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and rearranging (5), we get
@@
$\varphi_{n,k,\gamma,\lambda}(s)$ $=\mu_{n,k,\gamma,\lambda}(\ell+1)\left(\wp(s)-\wp(z)\right)\prod_{i=1}^{\ell}\wp(s)-\wp(z_{i})$
@@
Assume that $S_{r_{1}-1}\wp(z_{1},\ldots,z_{\ell}),S_{r_{2}-1}\wp(z_{1},\ldots,z_{\ell})\neq 0$. Then we have
$(-1)^{r_{1}}S_{r_{2}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{1})-(-1)^{r_{2}}S_{r_{1}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{2})=\left(S_{r_{1}}S_{r_{2}-1}-S_{r_{2}}S_{r_{1}-1}\right)\mu_{n,k,\gamma,\lambda}(\ell+1),$
where $S_{r}=S_{r}\wp(z_{1},\ldots,z_{\ell})$.
We note that $\mu_{n,k,\gamma,\lambda}(r)$ are computable in terms of $\gamma_{i},\lambda_{j}$ by Lemma 2.2 since the expansion of $\wp^{(i)}$ in terms of Weierstrass-$\wp$ is easily obtained by induction and differentiation. Applying Cramer’s rule to the system (4) yields
@@
where only $j$th column of the upper matrix has been replaced.
@@ 3 Applications And Discussion
We illustrate our method with the simplest case. We reobtain the classical proof of addition formula Theorem 1.1, as given in [9] and obtain two novel addition and duplication formulas in addition to new identities for elliptic invariants. Let $n_{1}=\gamma_{1}=1$, $(k_{1},k_{2})=(0,-2)$ and $\ell=2$. Then we obtain the system
@@
subject to the condition $\wp(z)\neq\wp(w)$. We then have $\psi(s)=\wp^{\prime}(s)-\lambda_{1}\wp(s)-\lambda_{2}$ and $\varphi(s)=\left(\wp^{\prime}(s)\right)^{2}-(\lambda_{1}\wp(s)+\lambda_{2})^{2}$. By equation (1), we calculate $\mu(r)$ as follows,
@@
Assume $z,w,z\pm w\notin\Lambda_{\Omega}$ and $z+w\not\equiv-z,-w\bmod\Lambda_{\Omega}$. Then we have
$\wp(z+w)=\frac{1}{4}\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}-\wp(z)-\wp(w).$ (8)
Furthermore, if $\wp(z)+\wp(w)\neq 0$, then we have
$\wp(z+w)=-\frac{\left(\wp^{\prime}(z)-\wp^{\prime}(w)\right)\left(\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)\right)}{2\left(\wp(z)+\wp(w)\right)\left(\wp(z)-\wp(w)\right)^{2}}-\frac{\wp(z)\wp(w)}{\wp(z)+\wp(w)}-\frac{g_{2}}{4\left(\wp(z)+\wp(w)\right)},$
if $\wp(z)\wp(w)\neq 0$, then we have
$\wp(z+w)=\frac{1}{4\wp(z)\wp(w)}\left(\frac{\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)}{\wp(z)-\wp(w)}\right)^{2}+\frac{g_{3}}{4\wp(z)\wp(w)}.$
@@
$\wp(2z)$ =′′$=-\frac{\wp^{\prime\prime}(z)\left(\wp^{\prime}(z)^{2}-\wp(z)\wp^{\prime\prime}(z)\right)}{4\wp(z)\wp^{\prime}(z)^{2}}-\frac{\wp(z)}{2}-\frac{g_{2}}{8\wp(z)}$
@@
which are valid for $\wp(z),\wp^{\prime}(z)\neq 0$, in other words $2z\notin\Lambda$. Above formula can be simplified further and written only in terms of $\wp(z)$ using $\left(\wp^{\prime}\right)^{2}=4\wp^{3}-g_{2}\wp-g_{3}$ and ′′$\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$. Indeed, we have
@@
Assume $z,w,z\pm w\notin\Lambda_{\Omega}$ and $z+w\not\equiv-z,-w\bmod\Lambda_{\Omega}$. If $\wp(z)+\wp(w)\neq 0$, then we have
@@
We now note several novel three-term addition formulas. Let $n_{1}=2$, $\gamma_{1}=1$, $(k_{1},k_{2},k_{3})=(1,0,-2)$ and $\ell=3$. We have the system
′′$\wp^{\prime\prime}(u)=\lambda_{1}\wp^{\prime}(u)+\lambda_{2}\wp(u)+\lambda_{3}$
′′$\wp^{\prime\prime}(v)=\lambda_{1}\wp^{\prime}(v)+\lambda_{2}\wp(v)+\lambda_{3}$
′′$\wp^{\prime\prime}(w)=\lambda_{1}\wp^{\prime}(w)+\lambda_{2}\wp(w)+\lambda_{3}.$
Using Frobenius-Stickelberger formula (2), condition (3) is equivalent to
@@
This directly translates to the assumptions that $u,v,w$ are distinct $\bmod\Lambda_{\Omega}$ and $u+v+w\notin\Lambda_{\Omega}$. The determinant in Theorem 2.1 unfortunately vanishes identically by the Frobenius-Stickelberger formula. We get that $\varphi(s)=\lambda_{1}^{2}\wp^{\prime}(s)^{2}-\left(\wp^{\prime\prime}(s)-\lambda_{2}\wp(s)-\lambda_{3}\right)^{2}$. Utilizing (1) and its corollary ′′$\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$, we calculate $\mu(r)$ as follows,
@@
Let $u,v,w$ be distict $\bmod\Lambda_{\Omega}$, $u,v,w,u+v+w\notin\Lambda_{\Omega}$ and $u+v+w\not\equiv-u,-v,-w\bmod\Lambda_{\Omega}$. Then we have
$\wp(u+v+w)=\frac{\lambda_{1}^{2}}{9}+\frac{\lambda_{2}}{3}-\wp(u)-\wp(v)-\wp(w).$ (9)
If $\wp(u)+\wp(v)+\wp(w)\neq 0$, then
$\wp(u+v+w)=\frac{\lambda_{2}^{2}-12\lambda_{3}-6g_{2}}{36(\wp(u)+\wp(v)+\wp(w))}-\frac{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}{\wp(u)+\wp(v)+\wp(w)}$
If $\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)\neq 0$, then
$\wp(u+v+w)=-\frac{\lambda_{1}^{2}g_{2}+2\lambda_{2}\lambda_{3}+\lambda_{2}g_{2}}{36(\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u))}-\frac{\wp(u)\wp(v)\wp(w)}{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}.$
If $\wp(u)\wp(v)\wp(w)\neq 0$, then
$\wp(u+v+w)=\frac{4\lambda_{1}^{2}g_{3}+4\lambda_{3}^{2}+4\lambda_{3}g_{2}+g_{2}^{2}}{144\wp(u)\wp(v)\wp(w)}.$
Now we take the limits $v,w\to u=z$ to obtain triplication formula for $\wp(z)$. After a lengthy computation of limits we obtain

### v2

@@ 1 Introduction
Let $\Omega=(2\omega_{1},2\omega_{2})\in\mathbb{C}^{2}$ be such that $\text{Im}(\omega_{2}/\omega_{1})>0$ and $\Lambda_{\Omega}=2\omega_{1}\mathbb{Z}+2\omega_{2}\mathbb{Z}$. The Weierstrass elliptic function $\wp$ is defined as
@@
The function $\wp$ satisfies the well-known algebraic differential equation
$\left(\wp^{\prime}(z)\right)^{2}=4\wp(z)^{3}-g_{2}\wp(z)-g_{3}=4(\wp(z)-e_{1})(\wp(z)-e_{2})(\wp(z)-e_{3}).$ (1.1)
@@
$\begin{vmatrix}1&\wp(z_{0})&\wp^{\prime}(z_{0})&\cdots&\wp^{(n-1)}(z_{0})\\ 1&\wp(z_{1})&\wp^{\prime}(z_{1})&\cdots&\wp^{(n-1)}(z_{1})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&\wp(z_{n})&\wp^{\prime}(z_{n})&\cdots&\wp^{(n-1)}(z_{n})\end{vmatrix}=(-1)^{\frac{n(n-1)}{2}}1!2!\cdots n!\frac{\sigma(z_{0}+z_{1}+\cdots+z_{n})}{\sigma(z_{0})^{n+1}\cdots\sigma(z_{n})^{n+1}}\prod_{0\leq i<j\leq n}\sigma(z_{i}-z_{j}),$ (1.2)
@@
Frobenius-Stickelberger identities may be used to obtain algebraic addition theorems for the function $\wp$ (for example, see [3, 4]). For more related determinant identities we refer the reader to [1, 2, 5, 6, 7]. The classical addition theorem for $\wp$ is given as follows.
@@
The addition formula for the $\wp$ function is crucial and in fact a characteristic property of the $\wp$ function, as shown in [8]. For recent advances in addition theorems for elliptic and abelian functions, see [9, 10]. Recent work in the research of addition formulas by these authors uses a more algebraic geometric approach and is applicable to the general case of higher genus curves. The authors have a long series of publications discussing such investigations. However, the addition formulas obtained are not explicit in the classical sense. For example, the general addition theorems obtained for the classical elliptic (genus 1) case make use of conjugate variables. Even in the Weierstrass case, calculation of conjugate variables reduce to solving the equation $\wp^{\prime}(u)=\wp^{\prime}(v)=\wp^{\prime}(w)$, which can not be easily done. The conjugate variables can be found explicitly only in the cases of special lattices which have particular symmetries, such as the equianharmonic and the lemniscatic case. Generally, one obtains Frobenius-Stickelberger like determinant identities. In this note, we discuss a recipe to obtain algebraic addition theorems for $\wp$, expressing $\wp(z_{1}+\cdots+z_{\ell})$ in terms of $\wp^{(j)}(z_{i})$ for predetermined $j$. The novelty of our approach is the immense flexibility in ones ability to choose parameters. One can choose which derivatives of the function $\wp$ and what constants can appear and modify them to obtain different addition formulas. Furthermore, our equations make no use of conjugate variables and only depend on the values of given variables $z_{1},\ldots,z_{\ell}$. Another thing to note is that we make no assumptions on the lattice.
@@ 2 Main Results
$\begin{vmatrix}\wp^{(k_{1})}(z_{1})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \wp^{(k_{1})}(z_{2})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \wp^{(k_{1})}(z_{\ell})&\wp^{(k_{2})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\end{vmatrix}\neq 0.$ (2.1)
@@
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{j})=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(z_{j})$ (2.2)
holds for every $j=1,\ldots,\ell$. Let us denote $z=z_{1}+\cdots+z_{\ell}$. Finally, assume that $z\notin\Lambda_{\Omega}$ and both of the equations $n_{a}=k_{b}$, $\gamma_{a}=\lambda_{b}$ do not hold at the same time. In order to obtain addition formulas, we shall prove that z-z is another solution of the equation (2.2). The assumptions above guarantee that the poles of the function $\wp$ are not encountered. We will see in the future that these assumptions may be relaxed.
We first prove the determinant form of the addition theorem.
@@
By our assumptions, $\psi_{n,k,\gamma,\lambda}$ is an elliptic function of order $\max(n_{a},k_{b})+2=\ell+1$ with a pole of order $\ell+1$ at $s=0$. Since the equation (2.2) holds for any $j=1,\ldots,\ell$, the function $\psi_{n,k,\gamma,\lambda}$ has $\ell$ distinct zeros, namely $z_{1},\ldots,z_{\ell}$. Let $w$ denote the $\ell+1$th zero. Using Abel’s theorem we obtain $w+\sum_{i=1}^{\ell}z_{i}\equiv 0\bmod\Lambda_{\Omega}$ and $w\equiv-z\bmod\Lambda_{\Omega}.$ Thus z-z is a root of $\psi_{n,k,\gamma,\lambda}$ and we get
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(-z).$ (2.3)
Putting equations (2.2) and (2.3) together gives
$\begin{pmatrix}\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\wp^{(k_{1})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\wp^{(k_{1})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\wp^{(k_{1})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)&\wp^{(k_{1})}(-z)&\cdots&\wp^{(k_{\ell})}(-z)\end{pmatrix}\begin{pmatrix}-1\\ \lambda_{1}\\ \vdots\\ \lambda_{\ell}\end{pmatrix}=-\begin{pmatrix}\psi_{n,k,\gamma,\lambda}(z_{1})\\ \psi_{n,k,\gamma,\lambda}(z_{2})\\ \vdots\\ \psi_{n,k,\gamma,\lambda}(z_{\ell})\\ \psi_{n,k,\gamma,\lambda}(-z)\end{pmatrix}=0.$ (2.4)
Since the vector on the left-hand side of (2.4) is non-zero, the matrix on the left-hand side is singular. Therefore we have
@@
We remark that since $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and $\wp$$\wp^{(i)}$ satisfy an algebraic relationship for every $i$, Theorem 2.1 is indeed an algebraic addition theorem.
@@
For all $n\in\mathbb{N}$, $\wp^{(2n)}\in\mathbb{Q}[g_{2},g_{2}][\wp]$ and $\wp^{(2n+1)}\in\wp^{\prime}\mathbb{Q}[g_{2},g_{2}][\wp]$. Furthermore $\wp^{(2n)}$ is a polynomial in $\wp$ of degree $n+1$ and $\wp^{(2n+1)}/\wp^{\prime}$ a polynomial in $\wp$ of degree $n$.
@@
The proof follows immediately after induction and equation (1.1). ∎
@@
$\varphi_{n,k,\gamma,\lambda}(s)=\left(\sum_{2\nmid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\nmid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}-\left(\sum_{2\mid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\mid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}$ (2.5)
@@
By Lemma 2.2 and equation (1.1), the first sum in (2.5) is in $\left(\wp^{\prime}\mathbb{C}[\wp]\right)^{2}$ and also in $\mathbb{C}[\wp]$. Similarly the second sum in (2.5) is a polynomial in $\mathbb{C}[\wp]$. Therefore $\varphi_{n,k,\gamma,\lambda}\in\mathbb{C}[\wp]$. Again by Lemma 2.2 and (1.1), the first sum is a polynomial in $\wp$ of degree at most $\max_{2\nmid n_{i},k_{j}}(n_{i},k_{j})+2$ and second sum is a polynomial in $\wp$ of degree at most $\max_{2\mid n_{i},k_{j}}(n_{i},k_{j})+2$. Thus $\varphi_{n,k,\gamma,\lambda}$ is a polynomial in $\wp$ of degree at most $\max_{i,j}(n_{i},k_{j})+2=\ell+1$. ∎
@@
Let $S_{r}(x_{1},\cdots x_{n})$, $r\leq n$ denote the elementary symmetric polynomials defined by
$S_{r}(x_{1},\ldots,x_{n})=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}x_{i_{1}}\ldots x_{i_{r}}.$
@@
Therefore $\wp(z_{1}),\ldots,\wp(z_{\ell})$ are distinct roots of $\varphi_{n,k,\gamma,\lambda}$. Using the fact that $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and rearranging (2.3), we get
@@
$\varphi_{n,k,\gamma,\lambda}(s)$ $=\mu_{n,k,\gamma,\lambda}(\ell+1)\left(\wp(s)-\wp(z)\right)\prod_{i=1}^{\ell}\left(\wp(s)-\wp(z_{i})\right)$
@@
Let us denote $S_{r}=S_{r}\wp(z_{1},\ldots,z_{\ell})$. Then we have
$(-1)^{r_{1}}S_{r_{2}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{1})-(-1)^{r_{2}}S_{r_{1}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{2})=\left(S_{r_{1}}S_{r_{2}-1}-S_{r_{2}}S_{r_{1}-1}\right)\mu_{n,k,\gamma,\lambda}(\ell+1).$
We note that $\mu_{n,k,\gamma,\lambda}(r)$ are computable in terms of $\gamma_{i},\lambda_{j}$ by Lemma 2.2 since the expansion of $\wp^{(i)}$ in terms of $\wp$ is easily obtained by induction and differentiation. Applying Cramer’s rule to the system (2.2) yields
@@
where only $j$th column of the upper matrix has been replaced. We now remark that, by meromorphic continuation of both sides in above equations, one might drop the initial assumptions. The formulas given hold for all values of the variables appearing in them.
@@ 3 Applications and Discussion
In this section we illustrate our method with two examples, reobtaining the classical addition formula and novel three-term addition formulas. The applications are not limited to these examples. Here we restrict ourselves to the case of two and three variables because of the difficulty in calculations.
Example 1.
Let us take $n_{1}=\gamma_{1}=1$, $(k_{1},k_{2})=(0,-2)$ and $\ell=2$. We reobtain the classical proof of addition formula Theorem 1.1, as given in [2] and obtain two novel addition and duplication formulas in addition to new identities for elliptic invariants. We obtain the system
@@
subject to the condition $\wp(z)\neq\wp(w)$. We then have $\psi(s)=\wp^{\prime}(s)-\lambda_{1}\wp(s)-\lambda_{2}$ and $\varphi(s)=\left(\wp^{\prime}(s)\right)^{2}-(\lambda_{1}\wp(s)+\lambda_{2})^{2}$. By equation (1.1), we calculate $\mu(r)$ as follows,
@@
We have the following addition formulae,
$\wp(z+w)$ $=\frac{1}{4}\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}-\wp(z)-\wp(w)$ (3.1)
$=-\frac{\left(\wp^{\prime}(z)-\wp^{\prime}(w)\right)\left(\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)\right)}{2\left(\wp(z)+\wp(w)\right)\left(\wp(z)-\wp(w)\right)^{2}}-\frac{\wp(z)\wp(w)}{\wp(z)+\wp(w)}-\frac{g_{2}}{4\left(\wp(z)+\wp(w)\right)}$
$=\frac{1}{4\wp(z)\wp(w)}\left(\frac{\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)}{\wp(z)-\wp(w)}\right)^{2}+\frac{g_{3}}{4\wp(z)\wp(w)}.$
@@
$\wp(2z)$ $=-\frac{\wp^{\prime\prime}(z)\left(\wp^{\prime}(z)^{2}-\wp(z)\wp^{\prime\prime}(z)\right)}{4\wp(z)\wp^{\prime}(z)^{2}}-\frac{\wp(z)}{2}-\frac{g_{2}}{8\wp(z)}$
@@
Above formula can be simplified further and written only in terms of $\wp(z)$ using $\left(\wp^{\prime}\right)^{2}=4\wp^{3}-g_{2}\wp-g_{3}$ and $\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$. Indeed, we have
@@
The following identity holds,
@@
Example 2.
Let us take $n_{1}=2$, $\gamma_{1}=1$, $(k_{1},k_{2},k_{3})=(1,0,-2)$ and $\ell=3$. We now note several novel three-term addition formulas. We have the system
$\wp^{\prime\prime}(u)=\lambda_{1}\wp^{\prime}(u)+\lambda_{2}\wp(u)+\lambda_{3}$
$\wp^{\prime\prime}(v)=\lambda_{1}\wp^{\prime}(v)+\lambda_{2}\wp(v)+\lambda_{3}$
$\wp^{\prime\prime}(w)=\lambda_{1}\wp^{\prime}(w)+\lambda_{2}\wp(w)+\lambda_{3}.$
Using Frobenius-Stickelberger formula (1.2), condition (2.1) is equivalent to
@@
This directly translates to the assumptions that $u,v,w$ are distinct $\bmod\Lambda_{\Omega}$ and $u+v+w\notin\Lambda_{\Omega}$. These conditions ensure that we do not encounter the poles of the $\wp$ function. The determinant in Theorem 2.1 unfortunately vanishes identically by the Frobenius-Stickelberger formula. We get that $\varphi(s)=\lambda_{1}^{2}\wp^{\prime}(s)^{2}-\left(\wp^{\prime\prime}(s)-\lambda_{2}\wp(s)-\lambda_{3}\right)^{2}$. Utilizing (1.1) and its corollary $\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$, we calculate $\mu(r)$ as follows,
@@
We have the following addition formulae,
$\wp(u+v+w)$ $=\frac{\lambda_{1}^{2}}{9}+\frac{\lambda_{2}}{3}-\wp(u)-\wp(v)-\wp(w)$ (3.2)
$=\frac{\lambda_{2}^{2}-12\lambda_{3}-6g_{2}}{36(\wp(u)+\wp(v)+\wp(w))}-\frac{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}{\wp(u)+\wp(v)+\wp(w)}$
$=-\frac{\lambda_{1}^{2}g_{2}+2\lambda_{2}\lambda_{3}+\lambda_{2}g_{2}}{36(\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u))}-\frac{\wp(u)\wp(v)\wp(w)}{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}$
$=\frac{4\lambda_{1}^{2}g_{3}+4\lambda_{3}^{2}+4\lambda_{3}g_{2}+g_{2}^{2}}{144\wp(u)\wp(v)\wp(w)}.$
Similarly, comparing above equations for different sides of $\wp(u+v+w)$ and rearranging gives us a variety of novel identities involving the invariants $g_{2}$ and $g_{3}$. Now we take the limits $v,w\to u=z$ to obtain triplication formula for $\wp(z)$. After a lengthy computation of limits we obtain